### v1

# Demo: Enabling Deep Reinforcement Learning Research for Energy Saving in Open RAN

Abstract
The growing performance demands and higher deployment densities of next-generation wireless systems emphasize the importance of adopting strategies to manage the energy efficiency of mobile networks. In this demo, we showcase a framework that enables research on Deep Reinforcement Learning (DRL) techniques for improving the energy efficiency of intelligent and programmable Open RAN systems. Using the open-source simulator ns-O-RAN and the reinforcement learning environment Gymnasium, the framework enables to train and evaluate DRL agents that dynamically control the activation and deactivation of cells in a 5G network. We show how to collect data for training and evaluate the impact of DRL on energy efficiency in a realistic 5G network scenario, including users’ mobility and handovers, a full protocol stack, and 3rd Generation Partnership Project (3GPP)-compliant channel models. The tool will be open-sourced upon acceptance of this paper and a tutorial for energy efficiency testing in ns-O-RAN.
      This paper has been accepted to IEEE CCNC 2025. If you wish to cite this work, please use the following reference:                         M. Bordin, A. Lacava, M. Polese, F. Cuomo, and T. Melodia,       “Demo: Enabling Deep Reinforcement Learning Research for Energy Saving in Open RAN,” in Proc. IEEE CCNC 2025.
I Introduction
The growing complexity of 5th generation (5G) Radio Access Network (RAN), driven by diverse use cases and rising data demands, has also led to increasing energy consumption in mobile networks [9678321]. This is driven by the use of larger bandwidths, an increasing number of RF chains and power amplifiers in the radio frontends, as well as dense, always-on, and inflexible deployments, which often negate improvements in the efficiency of individual components. [9678321, salem:tel-02500618, 7060678]. However, the emergence of Open RAN presents the potential for dynamic tuning and optimization of network systems [polese2023understanding]. This opens new avenues for reconfiguring the infrastructure to adapt to the fluctuating workloads of cellular networks, enabling resource scaling as needed and, consequently, offering opportunities to minimize energy consumption while maintaining performance requirements. Efficiently achieving this, however, remains an open challenge. Simultaneously, tools for conducting research and evaluation with realistic protocol stacks, dynamic control, and accurate channel models are still limited. In this demo, we introduce a new learning environment [^1] https://github.com/wineslab/ns-o-ran-gym specifically designed for energy-saving research in Open RAN systems made by extending the popular Deep Reinforcement Learning (DRL) framework, Gymnasium [10619796], and its integration with the open-source simulator ns-3 and the Open RAN ns-O-RAN library [lacava2023ns]. In our previous work [ESoran], we introduced energy-saving functionalities in ns-O-RAN, an open-source module part of the OpenRAN Gym suite for 5G RAN modeling, allowing for realistic xApp evaluation using 3rd Generation Partnership Project (3GPP)-based protocols [lacava2023ns]. This demo goes a step further by equipping the research community with a user-friendly tool to develop DRL solutions for optimizing energy savings while meeting Quality of Service (QoS) requirements. The Gymnasium environment provides an Application Programming Interface (API) to ns-O-RAN, enabling the testing of online DRL, where cell activation and deactivation serve as control actions. As part of the demonstration, we will present the integrated framework and show how to develop a DRL agent using Proximal Policy Optimization (PPO), along with heuristic baselines for comparison.
II Framework Design
Figure 1: System architecture
The demo leverages two main components. First, the ns-O-RAN module [lacava2023ns], a simulation environment based on ns-3, which models the RAN and its connectivity to the near-real-time RAN Intelligent Controller (RIC). Second, a control environment based on Gymnasium, a Python library for creating reinforcement learning environments [10619796]. This library expands the ns-O-RAN framework capabilities to support online DRL providing a standard an API interface for ns-O-RAN. The framework provides the NsOranEnv class, an abstract base class for creating environments compliant with the gym framework, specifically designed for ns-O-RAN simulations. It provides methods that another class must override to provide additional information to shape a specific network optimization problem. This demo presents EnergySavingEnv, a class implementing the Energy Saving use case. We extend the abstract methods of NsOranEnv according to the problem specifications. Action space. Each action corresponds to the decision to either activate or deactivate a specific cell based on network conditions and maximize the reward function. Since the environment involves NN cells, the action space includes $2^{N}$ possible actions, represented as an NN-bit binary list by the DRL model. _compute_action(), is a helper function that converts the agent’s action defined in Gymnasium into the format required by ns-O-RAN before triggering the ActionController class. Observation space. It consists of 12 cell-level Key Performance Measurements (KPMs) and one scenario-level KPM, for a total of $12\cdot N+1$ KPMs, where NN is the number of cells in the scenario. These KPMs are selected from the system’s available states based on their high correlation with the reward function. The method _get_obs() returns the observation state. Reward function. The reward is a weighted combination of a list of KPMs, incorporating throughput, the energy consumption of the connected base stations, the activation cost for a cell, and an exponential decay function. This decay ensures higher costs when action changes occur in quick succession, and lower costs when there is a longer gap between changes. _compute_reward(), is the method that returns the reward function based on the observation state; Finally, the function _fill_datalake_usecase() captures additional data from ns-O-RAN and stores it in the Datalake. However, since the Datalake uses the cell IMSI and timestamp as its primary key, and some specific Energy Saving KPMs are cell-centric, we do not store the extracted additional features in the Datalake, but we keep track of them within the custom environment as class variables.
Figure 2: Simulation scenario overview.
III Demonstration
The demonstration involves setting up an environment simulating a dense urban area based on the 3GPP UMi Street Canyon model [3gpp.38.913, 3gpp.38.901]. Fig. 2 shows a 5G NSA configuration with a central LTE eNB and Next Generation Node Base (gNB), surrounded by additional gNBs spaced 1700 meters apart. The near-real-time RIC, deployed at the RAN edge, manages the network via the E2 interface, collecting performance data and controlling both LTE and 5G cells (Fig. 1). The scenario includes 63 User Equipments (UEs) positioned uniformly with a random walk mobility model. Downlink traffic varies across UEs, simulating a mix of TCP and UDP applications. Specific cells’ RF frontends are deactivated to reduce RAN power consumption, as per the power-saving strategy outlined in [9678321]. We will show how to use the ns-O-RAN API to set up the simulation. Then, during the simulation, every 100 ms, the DRL model analyzes the KPMs collected by the KPM collector and dynamically activates or deactivates the Radio Frequency (RF) frontend of the gNB deployed in the scenario. We will display energy consumption and throughput performance on an interactive dashboard, enabling real-time monitoring and analysis. This interface facilitates a comprehensive comparison between the performance of the DRL model and traditional heuristic methods. We test the DRL agent introduced in [ESoran], designed with a reward function that models throughput, energy consumption, overage, and cell activation costs.

### v2

@@
 Cuomo, and T.
-Melodia,       “Demo: Enabling Deep Reinforcement Learning Research for Energy Saving in Open RAN,” in Proc.
-IEEE CCNC 2025.
+Melodia,                “Demo: Enabling Deep Reinforcement Learning Research for Energy Saving in Open RAN”, 2025 IEEE 22nd Consumer Communications & Networking Conference (CCNC), pp.1–2.
+doi:10.1109/CCNC54725.2025.10975928
 I Introduction